Corrections:

Let f(x) = x2 + 2x + 4. Which of the following statements is NOT true? A. f(x) has a maximum value B. The graph of f is not a line C. The graph of f has no x-intercepts. D. The graph of f has a y-intercept.

MusicGeek:

I think it is D, the graph of f has a y-intercept.

Corrections:

And could you explain why you think the answer is D?

YRJ8498:

it is a parabola opening up. so f(x) has a minimum thus A is not true

MusicGeek:

@yrj8498 wrote:
it is a parabola opening up. so f(x) has a minimum thus A is not true
wait, this dude is correct. i did my math wrong.

YRJ8498:

@musicgeek wrote:
@yrj8498 wrote:
it is a parabola opening up. so f(x) has a minimum thus A is not true
wait, this dude is correct. i did my math wrong.
glad you took it like a champ tho!

Corrections:

@yrj8498 wrote:
it is a parabola opening up. so f(x) has a minimum thus A is not true
Well, in the most nice way possible it is.

YRJ8498:

@corrections wrote:
@yrj8498 wrote:
it is a parabola opening up. so f(x) has a minimum thus A is not true
Well, in the most nice way possible it is.
what you mean?

Corrections:

Fx has a maximum value because after the variable, the numbers are the same on both sides hence, A is the most accurate choice.

surjithayer:

A is not true. \[f(x)=x^2+2x+4=x^2+2x+1+3=(x+1)^2+3\] when x=0,f(x)=4 so it has a y-intercept. when f(x)=0,\[(x+1)^2=-3\] \[(x+1)=\pm \sqrt{-3}\] Hence no x-intercept. it is a quadratic,it is not a st. line. minimum value of f(x)=3 and is attained at x=-1 so only A is not true.

YRJ8498:

@surjithayer wrote:
A is not true. \[f(x)=x^2+2x+4=x^2+2x+1+3=(x+1)^2+3\] when x=0,f(x)=4 so it has a y-intercept. when f(x)=0,\[(x+1)^2=-3\] \[(x+1)=\pm \sqrt{-3}\] Hence no x-intercept. it is a quadratic,it is not a st. line. minimum value of f(x)=3 and is attained at x=-1 so only A is not true.
Which means A is the Answer good job